OpenStudy (anonymous):

Eight subtracted from 4 times a number is equal to 4 less than 6 times the number. What is the number?

OpenStudy (anonymous):

The required expression is 8-4x=4-6x. Find x: -2x=2 and x=-1

OpenStudy (kamibug):

@Hoslos I think the expression should be arranged like this... 4x - 8 = 6x - 4 It says, "Eight subtracted from 4 times a number" not "4 times a number subtracted from 8" same goes for the "4 less than 6 times a number." Correct me if I'm wrong? :)

OpenStudy (anonymous):

I am not totally sure. You might be right.

OpenStudy (anonymous):

But our answers would be symmetric, right?

OpenStudy (kamibug):

@Hoslos They would be if it was -8+4x and -4+6x. The same way that 3-2 is not equal to 2-3 but it is equal to -2+3... :)

OpenStudy (anonymous):

Correct. I am getting more the essence. Yes, your expression must be the right one.

OpenStudy (kamibug):

@hurica1n 4x - 8 = 6x - 4 Then you'd combine like terms! :D Add 8 to both sides... 4x = 6x + 4 Subtract 6x from both sides... -2x = 4 Divide both sides by -2... x = 4/-2 x = ?

OpenStudy (anonymous):

-2